OpenStudy (anonymous):

What amount should an employee receive as bonus so that he would net 10 000 pesos after deducting 30 % for taxes?

OpenStudy (amistre64):

really?

OpenStudy (anonymous):

should be 10,000 x .30....that answer added to back to 10,000

OpenStudy (amistre64):

its still gonna be 10000/1.30 i beleive

OpenStudy (anonymous):

probably an easier way to solve that...but...use logic?

OpenStudy (anonymous):

after a 30% decrease you want 10,000 if you decrease a number by 30% that is the same as taking 70% of the number so you want 70% of a number equal to 10 000 write this in math replacing "of" by times "equals" by = 70% by .70 and "some number" by a variable i choose x \[.70\times x=10,000\]solve for x in one step

OpenStudy (anonymous):

@amistre and sonofa i think it is none of those, but i have not had coffee yet so i will not be money

OpenStudy (amistre64):

yeah, mines off :) it must be the pesos

OpenStudy (anonymous):

use the exchange rate to convert to euros, then it is easier

OpenStudy (amistre64):

A = P-P(.30) = P(1-.30) A/(1-.30) = P ....

OpenStudy (anonymous):

@sonofo if you add 30% back, it will not work because 30% of your reduced number is less that 30% of your original number. this is a common error. it is a division problem, not an addition one

OpenStudy (anonymous):

yea i shouldn't be doing math before my morning coffee...i know my brain doesn't work right yet...just being stubborn